SELF-BROADENING COEFFICIENTS FROM FTIR MEASUREMENTS OF H_{2}O

\underline{KAREN A. KEPPLER} and K. NARAHARI RAO, Department of Physics, The Ohio State University, Columbus, OH 43210-1106; Vl. G. TYUTEREV and S. N. MIKHAILENKO, GSMA U.A. CNRS D 1434, Faculté des Sciences, BP 347-51062 Reims Cédex France; G. Ch. MELLAU, S. KLEE, B. P. WINNEWISSER, and M. WINNEWISSER, Physikalisch-Chemisches Institut, Justus-Liebig-Universität Giessen, D-35392 Giessen, Germany.

Self-broadening parameters have been extracted for selected transitions in the ν_{1}, 2ν_{2} and ν_{3} bands of water vapor observed between 2500–4200 cm^{−1}. The H_{2}O spectrum at room temperature in natural isotopic abundance has been recorded in this region at pressure–broadening–limited resolution with a Bruker FTS 120HR and a White–type cell adjusted to provide pathlengths of 64.5–288.5m. Sample pressures used were 0.007–29.7mbar (0.005–22.3 Torr). Broadening coefficients determined for transitions in the ν_{2} band of water vapor in this study and earlier works^{a} are comparable. The results of the broadening determinations are presented and discussed.

aJ. -Y.Mandin, C. Camy-Peyret, and J. -M. Flaud, J. Quant. Spectrosco. Radiat. Transfer 26(6), 483–494 (1981).